Solve the problem. A man rode a bicycle for 12 miles and then hiked an additional 8 miles. The total time for the trip was 5 hours. If his rate when he was riding a bicycle was 10miles per hour faster than his rate walking, wat was each rate?
Answer by mananth(16946)

A man rode a bicycle for 12 miles and then hiked an additional 8 miles. The total time for the trip was 5 hours. If his rate when he was riding a bicycle was 10miles per hour faster than his rate walking, wat was each rate?
let his rate of walking be x mph
His rate of riding will be x + 10 mph
.
time taken for walking = 8/x hours
time taken for riding = 12/x+10
.
Time walking + time riding = total time
8/x + 12/ x+10 = 5 hours
LCM = x(x+10)
8(x+10) + 12x / x (x+10)= 5
8x +80 +12x = 5x^2 +50x
5x^2+30x-80=0
x^+6x-16 = 0
x^2+8x-2x-16=0
x(x+8)-2(x+8)=0
(x+8)(x-2)=0
x=2 which is positive
He walks at 2 mph
rides at 12 mph